What interest rate would you need in order for $2000 to double in 4 years if the amount is compounded monthly? Show formula used
Answer #1

We use the formula:
A=P(1+r/12)^12n
where
A=future value
P=present value
r=rate of interest
n=time period.

(2*2000)=2000*(1+r/12)^(12*4)

(2)^(1/48)=(1+r/12)

(1+r/12)=1.014545335

r=(1.014545335-1)*12

=17.45%(Approx).